Ordered: Levothyroxine 0.1mg po daily. Available: Tablets labeled 25 mcg. How many tablets will the nurse administer?

Ordered: 3,000 units of Heparin SQ. THe vial contains 5,000 units/mL. How many mL of heparin will you give?

Ordered: Naloxone (Narcan) O.5mg IV stat for respiratory depression. The vial contains 0.2 mg/mL. How many mL will you draw up in the syringe?
Answer #1

Ans) Ordered: Levothyroxine 0.1mg po daily. Available: Tablets labeled 25 mcg.

25 mcg/1000 = 0.025 mg

0.025 mg = 1 tab

0.1 mg = x

x = 1 × 0.1 / 0.025 = 4 tablets.

Ordered: 3,000 units of Heparin SQ. THe vial contains 5,000 units/mL

5000 units= 1 ml

3000 units = x

x = 3000/5000 = 0.6 ml.

Ordered: Naloxone (Narcan) O.5mg IV stat for respiratory depression. The vial contains 0.2 mg/mL.

0.2 mg = 1 ml

0.5 mg = x

x = 0.5/0.2 = 2.5 ml